The hypotenuse of a right triangle is 10cm. one of the legs is 2cm longer than the other leg.Find the area of the triangle
Answer by mananth(16949)

10^2-2^2= side^2
100-4 = side ^2
sqrt96=side
area of triangle = 1/2 base * height
= 1/2 * 2*sqrt96
= sqrt96 is the area of the triangle
= 9.79 square cms